OpenStudy (anonymous):

If the graph of y = -4x + 7 is perpendicular to the graph of y = mx + 3 , then what does m equal? -4 4 1/4 -1/4 Given A(9, 11), B(3, -4), C(6, 9), and D(-7, 7), then the midpoint of which line is located in Quadrant II? AB BC AC BD

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

if I call your slope with m, then the subsequent equation has to be checked: m*M=-1, where M=-4, so m=-1/4

OpenStudy (michele_laino):

oops sorry: m=1/4

OpenStudy (anonymous):

lol ok wb #2?

OpenStudy (michele_laino):

applying the formula for midpoint, given two points, we have: midpoint AB= (6,7/2) midpoint AC=(15/2,10), midpoint BC=(9/2,5/2) midpoint BD=(-2,3/2) 

OpenStudy (anonymous):

so is the answer BC?

OpenStudy (michele_laino):

your answer is midpoint of BD, namely the fourth option

OpenStudy (anonymous):

oh ok ty!

OpenStudy (michele_laino):

thank you!